OpenStudy (anonymous):

How many different groups of 35 cars can be formed from 40 cars?

Parth (parthkohli):

40C35.

OpenStudy (anonymous):

658008

OpenStudy (anonymous):

thanks

Parth (parthkohli):

\( \color{Black}{\Rightarrow \Large {40! \over (40 - 35)!35!}}\) \( \color{Black}{\Rightarrow \Large {40! \over 5!35!}}\) \( \color{Black}{\Rightarrow \Large {40 \times 39 \times 38 \times 37 \times 26 \over 5 \times 4 \times 3 \times 2}}\)

OpenStudy (anonymous):

thanks